In a random sample of students, 25% chose maths as their favourite subject. How many students in the school of 1200 students would you expect to choose maths as their favourite subject?
1
Expert's answer
2019-09-11T10:29:58-0400

No. of students choosing maths as their favourite subject


"= percentage\ of\ student \ chosing \ math\times total\ no. \ of\ students"


"=\frac{25}{100}\times 1200=300\ students"